OpenStudy (anonymous):

obtain the positive solution less than 360 degree 4sin^2x-3 = 0

OpenStudy (amistre64):

add 3, divide off the 4; and sqrt it all

OpenStudy (amistre64):

sin(x) = sqrt(3)/2 = 60 and 120 if i see it right

OpenStudy (anonymous):

yes thats the formula but when i put it on http://www.wolframalpha.com/ it says false TT

OpenStudy (amistre64):

false TT? im gonna have to try that :)

OpenStudy (anonymous):

sin(x) = sqrt(3)/2 x= 60 and 120 yes false haha

OpenStudy (amistre64):

its, 60 and 120 degrees ...

OpenStudy (anonymous):

1/cos^2(x)=4; cos(x)=1/2; x=+/-acos(1/2)+2*Pi*k; x=60°; 300°;

OpenStudy (anonymous):

but i know its 60 and 120 is the answer but theres something is missing i forgot

OpenStudy (amistre64):

right, for cos(60) and cos(300) = 1/2 1/2^2 = 1/4 1/(1/4) = 4

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

"positive solution" means that only positive values for x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

weee i get it know ^^ thx